how that the set of all 3 x 3 upper
triangular matrices with entries from R is a
vector space over R under usual addition
and scalar multiplication of matrices. Find a
basis of this vector space.
1
Expert's answer
2020-05-11T11:25:56-0400

We will verify all the axiom which qualify the set of all "3 \times 3" upper triangular matrix forms a vector space over "\mathbb{R}" ,

Let's define "M:=\bigg\{
\begin{pmatrix}
 a & b & c\\
 0& d&e\\
 0&0&f
\end{pmatrix} : \forall a,b,c,d,e,f \in \mathbb{R} \bigg\}" . Now, let for any "A , B ,C \in M" such that "A = \begin{pmatrix}
 a_1 & b_1 & c_1\\
 0& d_1&e_1\\
 0&0&f_1
\end{pmatrix}" , "B=\begin{pmatrix}
 a_2 & b_2& c_2\\
 0& d_2&e_2\\
 0&0&f_2
\end{pmatrix}" ,"C=\begin{pmatrix}
 a_3 & b_3 & c_3\\
 0& d_3&e_3\\
 0&0&f_3
\end{pmatrix}" . We will prove the first four axioms i.e "M" is an Abelian group under addition.



"1. A+B =\begin{pmatrix}
 a_1+a_2 & b_1+b_2 & c_1+c_2\\
 0& d_1+d_2&e_1+e_2\\
 0&0&f_1+f_2
\end{pmatrix}=B+A" "\in M"

"2.(A+B)+C = \begin{pmatrix}
 a_1+a_2+a_3 & b_1+b_2+ b_3& c_1+c_2+c_3\\
 0& d_1+d_2+d_3&e_1+e_2+e_3\\
 0&0&f_1+f_2+f_3
\end{pmatrix}=A+(B+C)"

"3.A+O =O+A=A" ,where "O" null matrix.

"4. A +(-A)=(-A)+A=O"


Now, we will show that multiplicative and distributive axioms holds true,



"5.O\cdot A = O \: (clearly, trivial\: multiplication)"

"6. for, 1 \in \mathbb{R}, 
1\cdot A =A"

"7. \forall \alpha, \beta \in \mathbb{R} \: and \: A \in M, \: \alpha(\beta \cdot A)=\alpha \begin{pmatrix}
 \beta a_1 & \beta b_1 & \beta c_1\\
 0& \beta d_1&\beta e_1\\
 0&0&\beta f_1
\end{pmatrix}"

"=\begin{pmatrix}
 \alpha \beta a_1 & \alpha\beta b_1 & \alpha\beta c_1\\
 0&\alpha \beta d_1&\alpha\beta e_1\\
 0&0&\alpha\beta f_1
\end{pmatrix}=(\alpha\cdot \beta)A"

"8." for any "\alpha \in \mathbb{R} \& A,B \in M" ,

"\alpha(A+B)= \begin{pmatrix}
 \alpha (a_1+a_2) & \alpha( b_1+b_2 )& \alpha( c_1+c_2)\\
 0& \alpha(d_1+d_2)& \alpha(e_1+e_2)\\
 0&0& \alpha(f_1+f_2)
\end{pmatrix}" "="  "\alpha A + \alpha B"


"9." for any "\alpha, \beta \in \mathbb{R}" and "A \in M" ,

"( \alpha + \beta)A =\begin{pmatrix}
 ( \alpha + \beta)a_1 & ( \alpha + \beta)b_1 &( \alpha + \beta) c_1\\
 0& ( \alpha + \beta)d_1&( \alpha + \beta)e_1\\
 0&0&( \alpha + \beta)f_1
\end{pmatrix}" "=\alpha A + \beta A"


Therefore, the set "M" satisfy all the axioms of vector space,hence "M" is a vector space.


One basis of "M" is follows,


"B=" "\bigg\{" "\begin{pmatrix}
 1 & 0 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}" ,"\begin{pmatrix}
 0 & 1 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}" ,"\begin{pmatrix}
 0 & 0 & 1\\
 0& 0&0\\
 0&0&0
\end{pmatrix}" ,"\begin{pmatrix}
 0 & 0 & 0\\
 0& 1&0\\
 0&0&0
\end{pmatrix}" ,"\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&1\\
 0&0&0
\end{pmatrix}" ,"\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&0\\
 0&0&1
\end{pmatrix}" "\bigg \}"


proof:

Let, for any "a,b,c,d,e,f \in \mathbb{R}" , "a\begin{pmatrix}
 1 & 0 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+b\begin{pmatrix}
 0 & 1 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+c\begin{pmatrix}
 0 & 0 & 1\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+"

"d\begin{pmatrix}
 0 & 0 & 0\\
 0& 1&0\\
 0&0&0
\end{pmatrix}+e\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&1\\
 0&0&0
\end{pmatrix}+f\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&0\\
 0&0&1
\end{pmatrix}=\begin{pmatrix}
 a & b & c\\
 0& d&e\\
 0&0&f
\end{pmatrix} \in M" .

Hence, "B" spans "M" .


Now, we will check the linear independence of "B" .


let for any "a,b,c,d,e,f \in \mathbb{R}," "a\begin{pmatrix}
 1 & 0 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+b\begin{pmatrix}
 0 & 1 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+c\begin{pmatrix}
 0 & 0 & 1\\
 0& 0&0\\
 0&0&0
\end{pmatrix}+"

"d\begin{pmatrix}
 0 & 0 & 0\\
 0& 1&0\\
 0&0&0
\end{pmatrix}+e\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&1\\
 0&0&0
\end{pmatrix}+f\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&0\\
 0&0&1
\end{pmatrix}=\begin{pmatrix}
 a & b & c\\
 0& d&e\\
 0&0&f
\end{pmatrix} =" "\begin{pmatrix}
 0 & 0 & 0\\
 0& 0&0\\
 0&0&0
\end{pmatrix}" ,

"\implies a=b=c=d=e=f=0" , Hence, "B" satisfy the linear independence.Thus "B" is a basis of "M" .

Therefore, we are done.